OpenStudy (anonymous):

(nRT^-2)^32/R^-2T^32 Is there an easier way to do this one?

OpenStudy (anonymous):

i had n32r32/t96

OpenStudy (anonymous):

but i think its wrong

OpenStudy (psymon):

\[\frac{ (nRT ^{-2})^{32} }{ R ^{-2}T ^{32} }\]?

OpenStudy (anonymous):

that's correct

OpenStudy (anonymous):

is it \[(nRT)^{-2}?\]

OpenStudy (anonymous):

it's (nRT^-2)32/r^-2t^32 jen

OpenStudy (psymon):

Well, let's distribute that 32 power on top and see what exactly is going on: \\[\frac{ n ^{32}R ^{32}T ^{-2*32} }{ R ^{-2}T ^{32} } \] Now you can see that the 32 on top multiplies in to each power. So that T part on top I'll simplify: \[\frac{ n ^{32}R ^{32}T ^{-64} }{ R ^{-2}T ^{32} } \]

OpenStudy (psymon):

So looks like you got one slight thing off. That should be R^34 on top.

OpenStudy (anonymous):

psymon good i got that correct so i'm glad about that, but somewhere after i messed up again

OpenStudy (psymon):

Well, you have a R^-2 on bottom that gets moved up and added into the R^32 on top, combining to become R^34.

OpenStudy (anonymous):

okay so i did what you said now my final answer looks like n^32r^34/t96 is this correct

OpenStudy (psymon):

Yep :3

OpenStudy (anonymous):

okay i think i'm getting it finally thanks psymon

OpenStudy (psymon):

Thats good ^_^ Yeah, the exponent rules are really important in higher math, so good to understand them now.

OpenStudy (anonymous):

yes